sin 5x * cos 3x = sin 9x * cos 7x. Find the general solution of the equation.

Grade:11

2 Answers

Aman Bansal
12 years ago

Dear Arnab,

Use Formula,

sumSines1.gif

and,

sumCosPlus.gif

to reduce the expression as sum terms and then it can be easily solved.

Best Of Luck...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

 

 

Fyu
7 years ago
   sin5x.cos3x=sin6x.cos2x                                             
   Multiplying both the sides by 2
   2sin5x.cos3x=2sin6x.cos2x                                        
   sin(5x+3x)+sin(5x-2x)=sin(6x+2x)+sin(6x-2x)
   sin8x+sin2x=sin8x+sin4x    
   sin2x-sin4x=0 
   2sin(2x-4x/2)cos(2x+4x/2)                                          
   -2sinxcos3x=0                        
    X=npi         Or 3x=(2n+1)pi/2 -> x=(2n+1)pi/6
       Hope this helps you!